OpenStudy (anonymous):

Select the equation of the line that passes through the point (–2, –1) and has slope 5 in point-slope form

OpenStudy (anonymous):

A. (y – 1) = 5(x – 2) B. (x + 2) = 5(y – 1) C. (x – 2) = 5(y + 1) D. (y + 1) = 5(x + 2)